# Strong Lensing Effect and Quasinormal Modes of Oscillations of Black Holes in $\boldsymbol{f(R,T)}$ Gravity Theory

### v1

@@ III Bending of Light in the Strong Field Limit
The above equation characterizes various possible photon orbits, which are graphically represented in Fig. 2 for $\omega=0,$ $1/3$ and $-2/3$. The figure depicts the effective radial potential corresponding to different impact parameter values such as $\zeta<\zeta_{p}$,  $\zeta=\zeta_{p}$ and $\zeta>\zeta_{p}$, which describe three different orbits of a photon. Here, $\zeta_{p}$ is the critical or minimum impact parameter defined at the photon sphere radius $r_{p}$ [117, 116]. A photon with impact parameter $\zeta<\zeta_{p}$ is captured by the strong field of the BHs, while one with $\zeta>\zeta_{p}$, is deflected by the BHs when reaches $r=r_{0}$ and with $\zeta=\zeta_{p}$, whirls the BHs for sometimes at photon sphere radius $r_{p}$ before leaving it [120]. The photon sphere radius is the largest positive root of the equation [121, 43, 65],
@@
$\frac{3M}{r}-\frac{4(\beta\omega+\pi(6\omega+2))r^{-\frac{8(\beta\omega+\pi(6\omega+2))}{16\pi-\beta(\omega-3)}}}{16\pi-\beta(\omega-3)}-r^{-\frac{8(\beta\omega+\pi(6\omega+2))}{16\pi-\beta(\omega-3)}}-1=0.$ (16)
Figure 3: Behaviors of photon sphere radius $r_{p}$ and critical impact parameter $\zeta_{p}$ for minimally coupled $f(R,T)$ gravity BHs as a function of model parameter $\beta$ for $\omega=0,$ $1/3$ and $-2/3$ respectively.
Due to the complexity of this equation the photon sphere radius $r_{p}$ for different classes of BHs are computed numerically for a range of $\beta$ values and a graphical representation of their dependence on the parameter $\beta$ is shown in Fig. 3. Furthermore, at the point of closest approach $r=r_{0}$ the impact parameter $\zeta_{0}$ can be expressed from the orbit equation (13) as $\zeta_{0}=\sqrt{C(r_{0})/A(r_{0})}$. This expression allows us to determine the critical impact parameter at $r_{0}=r_{p}$ as given below:
$\zeta_{p}=\sqrt{\frac{C(r_{p})}{A(r_{p})}}=\frac{r_{p}}{\sqrt{1-\frac{2M}{r_{p}}+\,r_{p}^{-\frac{8(\beta\omega+\pi(6\omega+2))}{16\pi-\beta(\omega-3)}}}}.$ (17)
@@ III.1 Deflection angle in minimally coupled $f(R,T)$ gravity spacetime
$=2-\frac{6M}{r_{0}}+\left[1+\frac{4(\beta\omega+\pi(6\omega+2))}{(16\pi-\beta(\omega-3))}\right]2\,r_{0}^{\!-\frac{8(\beta\omega+\pi(6\,\omega+2))}{16\pi-\beta(\omega-3)}},$ (28)
@@
$=\frac{6M}{r_{0}}-\left[1+\frac{8(\beta\omega+\pi(6\omega+2))}{16\pi-\beta(\omega-3)}+\frac{4(\beta\omega+\pi(6\omega+2))\left(\frac{8(\beta\omega+\pi(6\omega+2))}{16\pi-\beta(\omega-3)}+1\right)}{16\pi-\beta(\omega-3)}\right]r_{0}^{-\frac{8(\beta\omega+\pi(6\omega+2))}{16\pi-\beta(\omega-3)}}-1.$ (29)
@@
$=\bar{a}\,\log\Bigg{[}2+\frac{[4M\,(16\pi-\beta(\omega-3))-8(\beta\omega+\pi(6\omega+2))\left(\frac{8(\beta\omega+\pi(6\omega+2))}{16\pi-\beta(\omega-3)}+1\right)r_{p}^{-\frac{8(\beta\omega+\pi(6\omega+2))}{16\pi-\beta(\omega-3)}+1}]}{(-\frac{2M}{r_{p}}+r_{p}^{-\frac{8(\beta\omega+\pi(6\omega+2))}{16\pi-\beta(\omega-3)}}+1)(16\pi-\beta(\omega-3))r_{p}}\Bigg{]},$ (33)
[…]
@@
Figure 4: Behaviors of lensing coefficients $\bar{a}$ (left) and $\bar{b}$ (right) in minimally coupled $f(R,T)$ gravity BHs spacetime as a function of model parameter $\beta$ for $\omega=0$, $1/3$ and $-2/3$ cases.
@@
To examine the behavior of the lensing coefficients in our modified BHs spacetime, we analyze their variation with model parameter $\beta$ in Fig. 4 and list their values in Tables 1, 2 and 3. For the $\omega=0$ case, the coefficient $\bar{a}$ initially increases moderately, reaches a peak around $\beta\approx-7$, then decreases as $\beta$ grows. When $\omega=1/3$, $\bar{a}$ exhibits a continuous but slow growth up to $\beta\approx 3.5$ then decreases smoothly as $\beta$ increases. In contrast, the coefficient $\bar{b}$ shows opposite trends. For the $\omega=0$ case, $\bar{b}$ decreases initially, reaches a minimum near $\beta\approx-7$, then increases as $\beta$ approaches to 00. After that it follows a smooth declination for $\beta>0$, while for the $\omega=1/3$ case, $\bar{b}$ falls gradually up to $\beta\approx 3.5$ from the negative end and shows a slow growth beyond this value. On the other hand, for the $\omega=-2/3$ case, we observe significant variations in the lensing coefficients within the range of $-13.7088\leq\beta\leq-9.4381$. Similar to the cases of $\omega=0$ and $\omega=1/3$, coefficients $\bar{a}$ and $\bar{b}$ exhibit analogous behavior in this case also. However, for $\beta>-9.4381$, $\bar{a}$ goes almost as flat and $\bar{b}$ declines more rapidly than in the previous two cases. In fact, we observe contrasting behaviors of the lensing coefficients as a function of $\beta$ across these scenarios.
Table 1: Estimation of numerical values of lensing coefficients and corresponding impact parameter $\zeta_{p}$ for the $f(R,T)$ BHs with $\omega=0$ case. The impact parameter is listed as $\zeta_{p}/R_{s}$, where $R_{s}$ is the Schwarzschild radius. Here and in the following tables SBH represents the Schwarzschild BH.
  $\beta$  $\bar{a}$ $\bar{b}$   $\zeta_{p}/R_{s}$   $\beta$ $\bar{a}$  $\bar{b}$ $\zeta_{p}/R_{s}$
- 16.7551 1 -0.40064 2.59808 SBH 1 -0.40063 2.59808
- 13.7088 1.02185 -0.41639 2.58851 2 0.95031 -0.41402 1.24511
- 12 1.10901 -0.47797 2.50511 5 0.89987 -0.44105 1.18795
- 9.4381 1.33024 -0.70592 2.17591 7 1.05798 -0.40061 1.15985
- 8. 1.43031 -0.83576 1.93813 9 0.85786 -0.47588 1.13709
- 5 1.26380 -0.54176 1.56361 10 0.85003 -0.48367 1.12725
0 0.99998 -0.40061 1.29904 12 0.83655 -0.49808 1.11004
Table 2: Estimation of numerical values of lensing coefficients and corresponding impact parameter $\zeta_{p}$ for the $f(R,T)$ BHs with $\omega=1/3$ case. The impact parameter is listed as $\zeta_{p}/R_{s}$, where $R_{s}$ is the Schwarzschild radius.
  $\beta$  $\bar{a}$ $\bar{b}$   $\zeta_{p}/R_{s}$   $\beta$ $\bar{a}$  $\bar{b}$ $\zeta_{p}/R_{s}$
- 18.8495 1 -0.40064 2.59808 SBH 1 -0.40063 2.59808
- 13.7088 1.04275 -0.43049 2.57354 2 1.43156 -0.83678 1.93099
- 12 1.09001 -0.46364 2.52719 5 1.43338 -0.82465 1.84445
- 9.4381 1.17096 -0.53002 2.42349 7 1.42298 -0.79790 1.79658
- 8 1.21757 -0.57512 2.35584 9 1.40717 -0.76392 1.75533
- 5 1.30930 -0.67943 2.21143 10 1.39807 -0.74589 1.73686
0 1.41412 -0.81755 1.99982 12 1.37873 -0.70998 1.70366
Table 3: Estimation of numerical values of lensing coefficients and corresponding impact parameter $\zeta_{p}$ for the $f(R,T)$ BHs with $\omega=-2/3$ case. The impact parameter is listed as $\zeta_{p}/R_{s}$, where $R_{s}$ is the Schwarzschild radius.
  $\beta$  $\bar{a}$ $\bar{b}$   $\zeta_{p}/R_{s}$   $\beta$ $\bar{a}$  $\bar{b}$ $\zeta_{p}/R_{s}$
- 13.7088 0.99949 -0.40002 2.59808 SBH 1 -0.40063 2.59808
- 13 1.00487 -0.40449 2.59681 -9.4381 0.$7$0.71668 0.91971
- 12 1.36119 -0.74704 2.12046 -9.7 0.72388 -0.67901 0.94701
- 11.5 1.22224 -0.49753 1.52078 0 0.61479 -1.15139 0.68800
- 11 0.92935 -0.42329 1.22188 7 0.61204 -1.23563 1.09545
- 10.5 0.80914 -0.53162 1.07373 10 0.61197 -1.25831 1.11004
@@
Also we illustrate the variation in the deflection angle resulting from the bending of light in the strong field limit of minimally coupled $f(R,T)$ gravity BHs as given by equation (30). This variation is analyzed as a function of the impact parameter $\zeta$ for different values of model parameter $\beta$, employing the values of $\zeta_{p}$ and lensing coefficients $\bar{a}$ and $\bar{b}$ determined from equations (17), (31) and (32) respectively. To compute $\bar{a}$ we rely on equations (22) and (29), evaluated at $r_{0}=r_{p}$. Meanwhile, the coefficient $\bar{b}$ is calculated by estimating $b_{R}$ numerically and determining $b_{D}$ using equations (31) in (33). Thus, the behaviors of the strong limit deflection angle $\hat{\alpha}(\zeta)$ are depicted for varying values of the model parameter $\beta$, separately for the cases of $\omega=0,$ $1/3$ and $-2/3$ respectively in Fig. 5.
@@
Figure 5: Variations of the deflection angle with the impact parameter $\zeta$ for different values of the model parameter $\beta$ for $\omega=0,$ $1/3$ and $-2/3$ cases respectively, where the black dashed curves represent the Schwarzschild BH. The points on the horizontal axes represent the values of the impact parameter $\zeta=\zeta_{p}$ at which the deflection angle diverges. $M=1$ is taken for all the cases.
@@
From each graphical representation presented in this figure, it is evident that $\hat{\alpha}(\zeta)$ increases as impact parameter $\zeta$ decreases and shows divergence at $\zeta=\zeta_{p}$ as expected. When compared with the Schwarzschild BH, the deflection angle of a minimally coupled $f(R,T)$ gravity BH is found smaller than the Schwarzschild case in all three scenarios. The deflection angle decreases with increasing $\beta$, eventually becoming negative for larger impact parameter values. This negativity is more pronounced for $\beta>0$ whereas for large negative values of $\beta$ i.e., for smaller values of $\beta$ the negativity reduces and the deflection angle approaches to resemble the Schwarzschild behavior beyond some specific impact parameter values. Regarding the negative deflection angle, it is noteworthy that at high impact parameter values photons experience a repulsive interaction with the BH’s gravitational field. Additionally, in the $\omega=-2/3$ case, the deflection angle attains a more negative value compared to the $\omega=0$ and $\omega=1/3$ scenarios and the corresponding $\omega=-2/3$ plot reveals a remarkably negative deflection angle for $\beta>-9.4381$. Moreover, it is seen that for a specific value of $\beta$, a photon exhibits divergence at a larger $\zeta_{p}$ in the $\omega=1/3$ case relative to others. These observations suggest that the gravitational field of a minimally coupled $f(R,T)$ gravity BH surrounded by the radiation field is relatively stronger compared to when it is surrounded by a dust or quintessence field. In fact, the appearance of a negative deflection angle suggests valuable insights into the gravitational behavior of BHs. Several pioneering studies have also highlighted the phenomenon of negative deflection angles in BH spacetime [127, 128, 129, 130, 116].

### v2

@@ III Bending of Light in the Strong Field Limit
The above equation characterizes various possible photon orbits, which are graphically represented in Fig. 3 for $\omega=0,$ $1/3$ and 2/3-2/3. The figure depicts the effective radial potential corresponding to different impact parameter values such as $\zeta<\zeta_{p}$,  $\zeta=\zeta_{p}$ and $\zeta>\zeta_{p}$, which describe three different orbits of a photon. Here, $\zeta_{p}$ is the critical or minimum impact parameter defined at the photon sphere radius $r_{p}$ [129, 128], at a point where $dV_{eff}(r/dr=0$. A photon with impact parameter $\zeta<\zeta_{p}$ is captured by the strong field of the BHs, while one with $\zeta>\zeta_{p}$, is deflected by the BHs when reaches $r=r_{0}$ and with $\zeta=\zeta_{p}$, whirls the BHs for sometimes at photon sphere radius $r_{p}$ before leaving it [132]. The photon sphere radius is the largest positive root of the equation [133, 48, 70],
@@
$\frac{3M}{r}-\frac{4c_{2}(\beta\omega+\pi(6\omega+2))r^{-\frac{8(\beta\omega+\pi(6\omega+2))}{16\pi-\beta(\omega-3)}}}{16\pi-\beta(\omega-3)}-c_{2}\,r^{-\frac{8(\beta\omega+\pi(6\omega+2))}{16\pi-\beta(\omega-3)}}-1=0.$ (16)
Due to the complexity of this equation, the photon sphere radius $r_{p}$ for different classes of BHs are computed numerically for a range of values of $\beta$ and $c_{2}$, and graphical representations of their dependence on the parameters are shown in Fig. 4. Now, as an example, to specify the location of the photon sphere for specific values of $\beta$ and $c_{2}$, we illustrate the behavior of both sides of the equation $2M/r-1=(n/2+1)c_{2}\,r^{-n}-M/r$, obtained from Eq. (16) in Fig. 5 as functions of $r$. It is seen from the figure that the curve for the RHS of the above equation intersects the solid curve for 2M/r12M/r-1 at two points, one of them (red dot marked point in Fig. 5) corresponds to the peak of the effective potential $V_{eff}$ (see Fig. 3), giving the position of the photon sphere in the region beyond event horizon $r_{eh}$ (see Fig. 2).
Figure 4: Behaviors of photon sphere radius $r_{p}$ and critical impact parameter $\zeta_{p}$ for minimally coupled $f(R,T)$ gravity BHs as a function of model parameter $\beta$ and $c_{2}$ for $\omega=0,$ $1/3$ and 2/3-2/3 respectively.
Refer to caption     Refer to caption     Refer to caption
Figure 5: Behavior of y=2M/r1y=2M/r-1 (solid line) and $y=(n/2+1)c_{2}\,r^{-n}-M/r$ (black dashed line) as functions of $r$. These plots are presented to locate the photon sphere radius $r_{p}$ of $f(R,T)$ BHs for $\omega=0,\,1/3,\,-2/3$ considering $M=1$. The red dot mark in each plot is the intersecting point of the solid curve and the black dashed line corresponding to the effective potential peak specifying the location of the photon sphere radius.
Furthermore, at the point of closest approach $r=r_{0}$ the impact parameter $\zeta_{0}$ can be expressed from the orbit equation (13) as $\zeta_{0}=\sqrt{C(r_{0})/A(r_{0})}$. This expression allows us to determine the critical impact parameter at $r_{0}=r_{p}$ as given below:
$\zeta_{p}=\sqrt{\frac{C(r_{p})}{A(r_{p})}}=\frac{r_{p}}{\sqrt{1-\frac{2M}{r_{p}}+\,c_{2}r_{p}^{-\frac{8(\beta\omega+\pi(6\omega+2))}{16\pi-\beta(\omega-3)}}}}.$ (17)
@@ III.1 Deflection angle in minimally coupled $f(R,T)$ gravity spacetime
$=2-\frac{6M}{r_{0}}+\left[1+\frac{4(\beta\omega+\pi(6\omega+2))}{(16\pi-\beta(\omega-3))}\right]2\,c_{2}\,r_{0}^{\!-\frac{8(\beta\omega+\pi(6\,\omega+2))}{16\pi-\beta(\omega-3)}},$ (28)
@@
$=\frac{6M}{r_{0}}-\left[1+\frac{8\,(\beta\omega+\pi(6\omega+2))}{16\pi-\beta(\omega-3)}+\frac{4\,(\beta\omega+\pi(6\omega+2))\left(\frac{8(\beta\omega+\pi(6\omega+2))}{16\pi-\beta(\omega-3)}+1\right)}{16\pi-\beta(\omega-3)}\right]c_{2}\,r_{0}^{-\frac{8(\beta\omega+\pi(6\omega+2))}{16\pi-\beta(\omega-3)}}-1.$ (29)
@@
$=\bar{a}\,\log\Bigg{[}2+\frac{[4M\,(16\pi-\beta(\omega-3))-8c_{2}\,(\beta\omega+\pi(6\omega+2))\left(\frac{8(\beta\omega+\pi(6\omega+2))}{16\pi-\beta(\omega-3)}+1\right)r_{p}^{-\frac{8(\beta\omega+\pi(6\omega+2))}{16\pi-\beta(\omega-3)}+1}]}{(-\frac{2M}{r_{p}}+c_{2}\,r_{p}^{-\frac{8(\beta\omega+\pi(6\omega+2))}{16\pi-\beta(\omega-3)}}+1)(16\pi-\beta(\omega-3))r_{p}}\Bigg{]},$ (33)
$b_{R}$ $=I_{R}(r_{p})=\int_{0}^{1}\!\big{[}N(z,r_{p})\,g(z,r_{p})-N(0,r_{p})\,g_{0}(z,r_{p})\big{]}\,dz.$ (34)
@@
Figure 6: Behaviors of lensing coefficients $\bar{a}$ (upper row) and $\bar{b}$ (lower row) in minimally coupled $f(R,T)$ gravity BHs spacetime as functions of model parameters $\beta$ and $c_{2}$ for $\omega=0$, $1/3$ and 2/3-2/3 cases.
@@
To examine the behavior of the lensing coefficients in our modified BHs spacetime, we analyze their variation with model parameters $\beta$ and $c_{2}$ in Fig. 6 and list their values in Tables 1, 2 and 3. For the $\omega=0$ case, the coefficient $\bar{a}$ initially increases moderately, reaches a peak around $\beta\approx-8$, then decreases as $\beta$ grows. When $\omega=1/3$, $\bar{a}$ exhibits a continuous but slow growth up to $\beta\approx-4.7$ then decreases smoothly as $\beta$ increases. In contrast, the coefficient $\bar{b}$ shows opposite trends. For the $\omega=0$ case, $\bar{b}$ decreases initially, reaches a minimum near $\beta\approx-10$, then increases as $\beta$ approaches to 2-2. After that it follows a smooth declination for $\beta>0$, while for the $\omega=1/3$ case, $\bar{b}$ falls gradually up to $\beta\approx-7.5$ from the negative end and shows a slow growth beyond this value. On the other hand, for the $\omega=-2/3$ case, we observe significant variations in the lensing coefficients within the range of $13.5\leq\beta\leq-7.5$. Similar to the cases of $\omega=0$ and $\omega=1/3$, coefficients $\bar{a}$ and $\bar{b}$ exhibit analogous behavior in this case also. However, for $\beta>-7.5$, $\bar{a}$ goes almost as flat and $\bar{b}$ declines rapidly than in the previous two cases beyond $\beta=-11$. In fact, we observe contrasting behaviors of the lensing coefficients as a function of $\beta$ across these scenarios. Moreover, these coefficients show nearly identical behavior with respect to parameter $c_{2}$ for both $\omega=0$ and $\omega=1/3$ cases. It is observed that initially they decrease slowly and then undergo a sudden fall beyond $c_{2}=1$. In contrast, the coefficients exhibit opposite trends with the parameter $c_{2}$ in the $\omega=-2/3$ case as presented in Fig. 6.
@@
Also we illustrate the variation in the deflection angle resulting from the bending of light in the strong field limit of minimally coupled $f(R,T)$ gravity BHs as given by equation (30). This variation is analyzed as a function of the impact parameter $\zeta$ for different values of model parameters $\beta$ and $c_{2}$, employing the values of $\zeta_{p}$ and lensing coefficients $\bar{a}$ and $\bar{b}$ determined from equations (17), (31) and (32) respectively. To compute $\bar{a}$ we rely on equations (22) and (29), evaluated at $r_{0}=r_{p}$. Meanwhile, the coefficient $\bar{b}$ is calculated by estimating $b_{R}$ numerically and determining $b_{D}$ using equations (31) in (33). Thus, the behaviors of the strong limit deflection angle $\hat{\alpha}(\zeta)$ are depicted for varying values of the model parameters $\beta$ and $c_{2}$, separately for the cases of $\omega=0,$ $1/3$ and 2/3-2/3 respectively in Fig. 7.
@@
Refer to caption   Refer to caption   Refer to caption
@@
Figure 7: Variations of the deflection angle with the impact parameter $\zeta$ for different values of the model parameters $\beta$ and $c_{2}$ for $\omega=0,$ $1/3$ and 2/3-2/3 cases respectively, where the black dashed curves represent the Schwarzschild BH. The points on the horizontal axes represent the values of the impact parameter $\zeta=\zeta_{p}$ at which the deflection angle diverges. $M=1$ is taken for all the cases.
Table 1: Estimation of numerical values of lensing coefficients and corresponding impact parameter $\zeta_{p}$ with respect to different values of parameters $\beta$ and $c_{2}$ for $f(R,T)$ BHs with $\omega=0$ case. The impact parameter is listed as $\zeta_{p}/R_{s}$, where $R_{s}$ is the Schwarzschild radius.
$\beta$  $\bar{a}$ $\bar{b}$ $\zeta_{p}/R_{s}$ $\beta$ $\bar{a}$ $\bar{b}$ $\zeta_{p}/R_{s}$ $c_{2}$ $\bar{a}$ $\bar{b}$ $\zeta_{p}/R_{s}$
- 16.731 1.00000 -0.40064 2.59808 SBH 1.00000 -0.40063 2.59808 0 1 -0.40064 1.29904
- 14 1.00660 -0.40543 2.59560 2 0.98156 -0.40447 1.89158 0.2 0.93188 -0.40192 1.21509
- 12 1.04324 -0.42545 2.55485 5 0.95891 -0.41319 1.82442 0.4 0.90693 -0.40549 1.12510
- 9.5 1.07920 -0.43546 2.42575 7 0.94658 -0.41979 1.78873 0.5 0.89005 -0.40895 1.07718
- 8 1.08128 -0.42885 2.33060 10 0.93135 -0.42981 1.74508 0.8 0.79887 -0.43947 0.91595
0 0.99998 -0.40061 1.94855 12 0.92291 -0.43624 1.72106 0.9 0.73501 -0.47014 0.85301
Table 2: Estimation of numerical values of lensing coefficients and corresponding impact parameter $\zeta_{p}$ with respect to different values of parameters $\beta$ and $c_{2}$ for $f(R,T)$ BHs with $\omega=1/3$ case. The impact parameter is listed as $\zeta_{p}/R_{s}$, where $R_{s}$ is the Schwarzschild radius.
$\beta$ $\bar{a}$ $\bar{b}$ $\zeta_{p}/R_{s}$ $\beta$ $\bar{a}$ $\bar{b}$ $\zeta_{p}/R_{s}$ $c_{2}$ $\bar{a}$ $\bar{b}$ $\zeta_{p}/R_{s}$
- 18.8495 1 -0.40064 2.59808 SBH 1 -0.40063 2.59808 0 1 -0.40064 1.29904
- 14 1.01666 -0.41168 2.58877 2 1.08114 -0.42854 2.32791 0.2 0.92310 -0.40918 1.26873
- 12 1.03728 -0.42266 2.56476 5 1.07858 -0.42476 2.29274 0.4 0.88874 -0.42410 1.23564
- 9.5 1.05859 -0.43179 2.52108 7 1.07644 -0.42241 2.27229 0.5 0.86539 -0.43570 1.21778
- 8 1.06756 -0.43458 2.49219 9 1.07413 -0.42024 2.25391 0.8 0.74655 -0.51098 1.15675
0 1.08204 -0.43109 2.35480 12 1.07059 -0.41736 2.22966 0.9 0.67062 -0.569633 1.13271
Table 3: Estimation of numerical values of lensing coefficients and corresponding impact parameter $\zeta_{p}$ with respect to different values of parameters $\beta$ and $c_{2}$ for $f(R,T)$ BHs with $\omega=-2/3$ case. The impact parameter is listed as $\zeta_{p}/R_{s}$, where $R_{s}$ is the Schwarzschild radius.
$\beta$ $\bar{a}$ $\bar{b}$ $\zeta_{p}/R_{s}$ $\beta$ $\bar{a}$ $\bar{b}$ $\zeta_{p}/R_{s}$ $c_{2}$ $\bar{a}$ $\bar{b}$ $\zeta_{p}/R_{s}$
- 13.7088 1.00004 -0.40064 2.59808 SBH 1 -0.40063 2.59808 0 1 -0.40064 2.59808
- 13 1.00239 -0.40252 2.59744 -9.5 0.82067 -0.57138 1.42679 0.2 1.08472 -0.68839 1.61999
- 12 1.08116 -0.43441 2.40080 -9 0.79594 -0.62411 1.35108 0.4 1.13480 -0.82506 1.22075
- 11.5 1.05033 -0.40639 2.12771 0 0.76393 -1.00503 1.00021 0.5 1.14925 -0.87285 1.09440
- 11 0.97273 -0.40731 1.86521 7 0.70173 -1.09473 0.94797 0.8 1.17446 -0.97636 0.84669
- 10 0.85517 -0.51272 1.52817 10 0.70107 -1.11889 0.93542 0.9 1.17947 -1.00255 0.74103
@@
From each graphical representation presented in this figure, it is evident that $\hat{\alpha}(\zeta)$ increases as impact parameter $\zeta$ decreases and shows divergence at $\zeta=\zeta_{p}$ as expected. When compared with the Schwarzschild BH, the deflection angle of a minimally coupled $f(R,T)$ gravity BH is found smaller than the Schwarzschild case in all the scenarios. The deflection angle decreases with increasing model parameters $\beta$ and $c_{2}$, eventually becoming negative for larger impact parameter values. This negativity is more pronounced for $\beta>0$ and the higher values of $c_{2}$, whereas for smaller values of $\beta$ and $c_{2}$ the negativity reduces and the deflection angle approaches to resemble the Schwarzschild behavior beyond some specific impact parameter values. Regarding the negative deflection angle, it is noteworthy that at high impact parameter values photons experience a repulsive interaction with the BH’s gravitational field. Additionally, in the $\omega=-2/3$ case, the deflection angle attains a more negative value compared to the $\omega=0$ and $\omega=1/3$ scenarios. Moreover, it is seen that for a specific value of $\beta$, and also $c_{2}$, a photon exhibits divergence at a larger $\zeta_{p}$ in the $\omega=1/3$ case relative to others. These observations suggest that the gravitational field of a minimally coupled $f(R,T)$ gravity BH surrounded by the radiation field is relatively stronger compared to when it is surrounded by a dust or quintessence field. In fact, the appearance of a negative deflection angle suggests valuable insights into the gravitational behavior of BHs. Several pioneering studies have also highlighted the phenomenon of negative deflection angles in BH spacetime [139, 140, 141, 142, 128].